# Doppler Pulse Amplification

### v1

@@ IV System Demonstration
Figure 3 demonstrates the space-time evolution, time-domain waveforms and frequency spectra of a space-time wedge implementation (Fig. 2) of the DoPA system (Fig. 1b), governed governed by the Doppler [Eq. (1)] and width [Eq. (2)] scaling relations. Initially, the incident pulse undergoes multiple comoving space-time reflections within the opening wedge, resulting in successive temporal stretching (Fig. 3b) and Doppler downshifting (Fig. 3c). Then, the expanded pulse is amplified, at the time indicated by the dashed lines in Fig. 3a and Fig. 3b. Finally, the interfaces begin moving toward each other (closing wedge), causing the pulse to undergo multiple contramoving space-time reflections, leading to successive temporal compression and Doppler upshifting. Ultimately, the pulse exits the system with its original duration and shape, but with significantly higher peak power. Notably, the amplifier gain, $G_{\text{a}}$, in Fig. 3 is only two, due to space constraints, though in practice it can reach $G_{\text{a}}=10^{3}-10^{4}$.
@@ V Device Size Reduction
Figure 4 depicts the space-time evolution, time-domain waveforms and frequency spectra of the space-time Fresnel implementation of the DoPA system. The pulse experiences the same sequence of Doppler-induced shifts and compansion as the wedge implementation (Fig. 3), as evidenced by the similar waveform and spectral evolution. The key difference lies in the oscillatory motion of the interfaces and, hence, the shorter spacings between the temporal pulses, which allows the device length to be significantly reduced. Note that precise synchronization between the two moving boundaries is critical to avoid spurious reflections from the sharp discontinuities. If not carefully managed, these interactions may degrade pulse quality and reduce overall efficiency.
@@
The contrast in device size between the wedge and Fresnel configurations is evident when comparing Fig. 3 and Fig. 4. To further illustrate this, Fig. 5 shows the case when $l_{\text{f}}=l_{\text{w}}$. In this scenario, the Fresnel profile allows for an additional expansion-compression pair, further reducing the peak power of the pulse. This enables a larger amplification gain—four in this case—before reaching the saturation threshold of the amplifier.
@@
Figure 5: DoPA using a space-time Fresnel structure with the same device size as in Fig. 3 ($l_{\text{f}}=l_{\text{w}}$). (a) Space-time diagram. (b) Time-domain waveforms, with the dashed line indicating the amplifying time. (c) Corresponding spectrum. Note that the dynamic ranges in (b) and (c) is twice those in Figs. 3 and 4.
VI Conclusions
@@
We introduced Doppler pulse amplification (DoPA), a novel approach that harnesses space-time modulation interfaces to amplify ultrashort pulses via Doppler frequency shifts, complementing established techniques like chirped pulse amplification (CPA) and divided-pulse amplification (DPA). We analyzed the system with a space-time wedge implementation, demonstrating how multiple space-time scattering events and associated Doppler frequency shifts can be utilized to achieve compansion, while also mitigating the gain narrowing problem inherent in CPA. Furthermore, we explored a space-time Fresnel architecture approach, which reduces system size without sacrificing amplification efficiency, making DoPA suitable for amplification of both picosecond and femtosecond pulses. This compact design makes DoPA particularly suitable for applications requiring high-power ultrashort pulses, especially where system size and complexity are critical factors.
Experimental realization of DoPA is potentially feasible across all frequency regimes. In the microwave regime, it can be realized by loading an artificial transmission line with varactor diodes controlled by a field-programmable gate array (FPGA). By sequentially switching these diodes in a specific pattern, it is possible to create two moving interfaces on the transmission line corresponding to the realization of space-time wedges. In optics, DoPA could be implemented using the concept of relativistic flying mirrors in plasma [40, 41], where high-intensity laser pulses ionize and accelerate electrons to create reflective plasma boundaries. These moving interfaces induce Doppler shifts and could be synchronized to emulate the space-time wedge or Fresnel configurations.

### v3

@@ IV System Demonstration
Figure 3 demonstrates the space-time evolution, time-domain waveforms and frequency spectra of a space-time wedge implementation (Fig. 2) of the DoPA system (Fig. 1b), governed by the Doppler [Eq. (1)] and temporal width [Eq. (2)] scaling relations. Initially, the incident pulse undergoes multiple comoving space-time reflections within the opening wedge, resulting in successive temporal stretching (Fig. 3b) and Doppler downshifting (Fig. 3c). Then, the expanded pulse is amplified, at the time indicated by the dashed lines in Fig. 3a and Fig. 3b. Finally, the interfaces begin moving toward each other (closing wedge), causing the pulse to undergo multiple contramoving space-time reflections, leading to successive temporal compression and Doppler upshifting. Ultimately, the pulse exits the system with its original duration and shape, but with significantly higher peak power. Notably, the amplifier gain, $G$, in Fig. 3 is only two, due to space constraints, though in practice it can reach $G=10^{3}-10^{4}$.
@@ V Device Size Reduction
Figure 4 depicts the space-time evolution, time-domain waveforms and frequency spectra of the space-time Fresnel implementation of the DoPA system (Fig. 1b). The pulse experiences the same sequence of Doppler-induced shifts and compansion as the wedge implementation (Fig. 3), as evidenced by the similar waveform and spectral evolution. The key difference lies in the oscillatory motion of the interfaces and, hence, the shorter spacings between the temporal pulses, which allows the device length to be significantly reduced. Note that precise synchronization between the two moving boundaries is critical to avoid spurious reflections from the sharp discontinuities. If not carefully managed, these interactions may degrade pulse quality and reduce overall efficiency.
@@
The contrast in device size between the wedge and Fresnel configurations is evident when comparing Figs. 3 and 4. To further illustrate this, Fig. 5 shows the case when $l_{\text{f}}=l_{\text{w}}$. In this scenario, the Fresnel profile allows for an additional expansion-compression pair, further reducing the peak power of the pulse. This enables a larger amplification gain—four in this case—before reaching the saturation threshold of the amplifier.
@@
Figure 5: DoPA using a space-time Fresnel structure with the same device size as in Fig. 3 ($l_{\text{f}}=l_{\text{w}}$). (a) Space-time diagram. (b) Time-domain waveforms, with the dashed line indicating the amplifying time. (c) Corresponding spectrum. Note that the dynamic ranges in (b) and (c) are twice those in Figs. 3 and 4.
VI Experimental Outlook
The DoPA system hinges on the ability to generate interfaces moving at relatively high speeds, posing significant experimental challenges. Although no direct method currently exists for creating such interfaces, recent advances in time-varying media [41, 42, 43, 44, 45, 46] provide a foundation that could be extended to space-time varying systems. Here, we outline how these existing platforms might be adapted to demonstrate a proof-of-concept realization of DoPA.
In the microwave regime, pure time interfaces have been realized by temporally modulating the effective impedance of artificial transmission lines, controlled via high-speed electronic components, such as electronic switches [41, 42] or high-speed photodiodes [43]. Extending this concept to space-time systems requires sequential, spatially coordinated switching to the ground that emulates a moving perfect electric conducting (PEC) interface, demanding precise spatial control of the switching signals. One strategy is to drive the switches using a field-programmable gate array (FPGA), allowing for precise tunable delays and sequencing. By synchronizing the activation of switches in a controlled pattern, it is possible to form two contramoving interfaces along the line, effectively realizing space-time wedges. Microwave implementations employing corporate feed networks with tailored optical delays offer another route. The key DoPA parameter is the normalized modulation velocity, $\beta=\Delta x/\left(c\Delta t\right)$, where $\Delta x$ is the distance between adjacent switching elements and $\Delta t$ is the time delay between their activation. To emulate sharp moving interfaces corresponding to the wedges, two conditions must be met. First, the local rise time at each site must remain shorter than $\Delta t$, ensuring that the wave perceives a distinct discontinuity. Second, $\Delta x$ must be small compared to the wavelength of the incoming pulse to ensure a smooth modulation profile. For instance, using parameters from [42]: $\Delta x=0.208$ m and $\Delta t=4$ ns yield $\beta=0.17$. With $\tau_{0}=0.5$ ns and $a=10^{2}$, the Fresnel device length [Eq. (4)] equals $l_{\text{f}}\approx 5.5$ m, which is smaller than the 6.246.246.24 m transmission line used in [41, 42].
Experiments in the optical regime have also demonstrated temporal discontinuities using pump-probe setups [44, 45, 46], which induce refractive index changes via intense pump pulses in nonlinear media. The main challenge lies in achieving index variations that are both sufficiently large and fast. Transparent conducting oxides, especially in the epsilon-near-zero regime, have emerged as promising candidates, capable of refractive index modulations approaching unity on femtosecond timescales. Space-time modulated configurations can be achieved by tilting the pump beam with respect to the probe beam [34]. Due to the oblique geometry, the pump-induced refractive index change sweeps across the medium, effectively forming a moving interface. Using two contratilted pump beams allows the creation of a space–time wedge structure. For instance, if the angle between the probe and pump pulse is 60°60$\degree$60 ° and the pump is traveling at the speed of light, the normalized modulation velocity is given by $\beta=1/\tan 60\degree=1/\sqrt{3}$. Next, taking $a=10^{3}$ and $\tau_{0}=10$ fs yields a Fresnel length $l_{\text{f}}\approx 3.5$ mm.
VII Conclusions
@@
We introduced Doppler pulse amplification (DoPA), a novel approach that harnesses space-time modulation interfaces to amplify ultrashort pulses via Doppler frequency shifts, complementing established techniques like chirped pulse amplification (CPA) and divided-pulse amplification (DPA). We analyzed the system with a space-time wedge implementation, demonstrating how multiple space-time scattering events and associated Doppler frequency shifts can be utilized to achieve compansion, while also mitigating the gain narrowing problem inherent in CPA. Furthermore, we explored a space-time Fresnel architecture approach, which reduces system size without sacrificing amplification efficiency, making DoPA suitable for amplification of ultrashort pulses. This compact design makes DoPA particularly suitable for applications requiring high-power ultrashort pulses, especially where system size and complexity are critical factors. Finally, we outlined potential experimental realizations of the DoPA system in both the microwave and optical regimes.